# SDP bounds on quantum codes

### v1

@@ Abstract.
This paper provides a semidefinite programming hierarchy based on state polynomial optimization to determine the existence of quantum codes with given parameters. The hierarchy is complete, in the sense that if a $(\!(n,K,\delta)\!)_{2}$ code does not exist then a level of the hierarchy is infeasible. It is not limited to stabilizer codes and thus applicable generally. While it is formally dimension-free, we restrict it to qubit codes through quasi-Clifford algebras. We derive the quantum analog of a range of classical results: first, from an intermediate level a Lovász bound for self-dual quantum codes is recovered. Second, a symmetrization of a minor variation of this Lovász bound recovers the quantum Delsarte bound. Third, a symmetry reduction using the Terwilliger algebra leads to semidefinite programming bounds of size $O(n^{4})$. With this we give an alternative proof that there is no $(\!(7,1,4)\!)_{2}$ quantum code, and show that $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$ codes do not exist.
GM and FH were supported by the Foundation for Polish Science through TEAM-NET (POIR.04.04.00-00-17C1/18-00). AN was supported by QLCI. We thank Christine Bachoc, Simeon Ball, Bram Bekker, Marcus Grassl, Carlos de Gois, Otfried Gühne, Kiara Hansenne, Moisés Bermejo Morán, Claudio Procesi, René Schwonnek, Andreas Winter, and Zhen-Peng Xu for fruitful discussions and helpful comments.
@@ 1. Introduction
Coding theory studies how to send messages error-free through a noisy channel. Typically, a classical message consists of a bit string of length $k$, with the noisy classical channel inducing errors that may flip some bits in the message. A error-correcting code allows one to encode the message with redundant information, so that it can be recovered after being sent through the channel. A paradigmatic example is the repetition code of block length $n$ where the message bit is repeated $n$ times, e.g., the bits 00 and $1$ are encoded as $000$ and $111$ respectively when $n=3$.
@@
A quantum message is represented by a positive semidefinite matrix of trace one acting on the $K$-dimensional Hilbert space $\mathds{C}^{K}$. That is, $\varrho\in L(\mathds{C}^{K})$ with $\varrho\succeq 0$ and $\operatorname{tr}(\varrho)=1$. A noisy quantum channel, represented in its Kraus decomposition by $\mathcal{N}(\varrho)=\sum_{E\in\mathcal{N}}E\varrho E^{\dagger}$ with $\sum_{E\in\mathcal{N}}E^{\dagger}E\preceq\mathds{1}$, may not only induce flip errors but also errors in the phases of the state. The repetition code is no longer valid for quantum codes due to the no-cloning theorem [73], and therefore more involved strategies have to be used for the error-correction of quantum messages. A quantum error-correcting code encodes a $K$-dimensional Hilbert space to a subspace of the $n$-qubit Hilbert space, such that the action of a noisy quantum channel can be corrected. The Knill-Laflamme conditions provide necessary and sufficient conditions for this to be possible, and can be formulated in terms of the projector $\Pi$ onto the code subspace: if $\Pi E^{\dagger}F\Pi=c_{EF}\Pi$ for all errors $E,F$ appearing in the channel $\mathcal{N}$, then there exists a recovery map $\mathcal{R}$ such that $\mathcal{R}(\mathcal{N}(\varrho))\propto\varrho$ [35]. The size $K$ of the quantum code then equals the rank of $\Pi$.
@@
One of the fundamental problems in coding theory is to identify the maximum code size $K$ for a given block length $n$ and minimum distance $\delta$. This problem is usually approached by defining weight enumerators [13, 67]: for classical linear codes $C$, these count the number of codewords of given Hamming weight, $A_{j}(C)=|\{c\in C\,:\,\operatorname{wt}(c)=j\}|$, where $\operatorname{wt}(c)$ is the number of non-zero positions in the bit string $c$. In turn, the quantum weight enumerators $A_{j}(\Pi)$ decompose the Hilbert-Schmidt norm $\operatorname{tr}(\Pi^{\dagger}\Pi)$ of the projector $\Pi$ into contributions according to the support of different quantum correlations [28]. One has $A_{j}=\sum_{E\in\mathcal{E}_{n}\,:\,\operatorname{wt}(E)=j}=\operatorname{tr}(\Pi E^{\dagger})\operatorname{tr}(\Pi E)$, where $\mathcal{E}_{n}$ is the $n$-qubit Pauli basis and $\operatorname{wt}(E)$ the number of coordinates in which $E$ acts nontrivially; whereas the dual enumerator is given by $B_{j}=\sum_{E\in\mathcal{E}_{n}\,:\,\operatorname{wt}(E)=j}=\operatorname{tr}(\Pi E^{\dagger}\Pi E)$. For stabilizer and codeword-stabilized quantum codes, the weight enumerators have combinatorial interpretations similar to their classical counterparts [47, 37].
The introduction of weight enumerators allows for efficiently computable linear and semidefinite programming bounds on the size of classical codes. To understand the origin of these classical code bounds, consider the Lovász number $\vartheta$ of the Hamming graph. A subset of vertices is termed independent, if no two of its vertices are adjacent. The Lovász number $\vartheta$ then bounds the independence number of a graph, that is the largest cardinality of an independent set, as $\alpha(G)\leq\vartheta(G)$. While the computation of $\alpha$ is NP-hard [32], $\vartheta$ can be expressed as semidefinite program [41]. A suitable symmetrization of the semidefinite program formulation of the Lovász theta number then yields the classical Delsarte bound, namely the condition that both the primal and dual enumerators of a code are non-negative. This result relies on the fact that given the indexing vector of a code $\chi$, the matrix $\chi\chi^{T}$ is positive semidefinite and has zero entries when the row and column codewords have a distance in $11,\dots,\delta-$ 1 [22].
We show that for self-dual quantum codes a quantum Delsarte bound arises in a similar fashion [Result B, Section 7], with the added constraint that the matrix also has zeros in rows and columns whose “codewords”, that is Pauli-strings, anticommute. As a consequence, it relates to a Lovász number for a graph with loops, and thus fits into a Lovász framework recently developed for hypergraphs [10]. Finally, following the strategy in Ref.  [23] we give efficiently computable semidefinite programming bounds on the size of quantum codes [Result C, Section 8]. Similar to the classical case, these scale as $O(n^{4})$, where $n$ is the block length, and build on the block-diagonalization of the Terwilliger algebra of the non-binary Hamming scheme. With this, we give an an alternative proof that there is no $(\!(7,1,4)\!)_{2}$ quantum code 111 Note that the original proof in Ref. [30] does not make use of mathematical optimization., as well as prove the non-existence of quantum codes with parameters $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$ [Section 9]. Thus the method is strictly stronger than the “enhanced” quantum linear programming bounds that include the shadow bounds by Rains [58, 57], which do not recover this result.
@@
Our second main result is that these bounds are part of a complete semidefinite programming (SDP) hierarchy [Result A, Section 4]. If a code with given parameters $(\!(n,K,\delta)\!)_{2}$ does not exist, then some level of the hierarchy will be infeasible. This hierarchy is based on the state polynomial optimization framework, which provides a sequence of outer approximations to the set of non-linear quantum correlations. Approximating $A_{j}$ one can also recover $B_{j}$ through the quantum MacWilliams identity, and impose the constraints $KB_{i}=A_{i}$ for $0\leq i<\delta$ based on the Knill-Laflamme conditions for quantum error correction. While this machinery is formally dimension-free, a characterization of quasi-Clifford algebras by Gastineau-Hills [21] pulls this result back to $n$-qubit systems.
@@
Overall, our approach is applicable to both stabilizer and non-stabilizer codes. It is based on moment matrices and thus offers high flexibility on the constraints to be included, so that small relaxations can be constructed. Lastly, because our formalism is dimension-free, it can also target scenarios where the Hilbert space is unknown.
1.1. Related work
Lai et al. [37] have introduced an SDP relaxation in the context of entanglement-assisted codeword stabilized codes. When restricted to the case with no pre-shared entanglement, their bounds showed no advantage over the LP bounds. Berta et al. [6] presented a SDP hierarchy to determine the existence of a quantum code based on the symmetric extension hierarchy [15]. While this hierarchy is also complete, it does not easily allow for numerically tractable relaxations. In the classical case, a complete hierarchy for the existence of codes can be formulated using $0-1$ programming for maximum independent sets [42, 38].
@@ 2. Contributions
Our aim is to determine the existence of quantum codes with block-length $n$, code distance $\delta$, and the dimension of the code space $K$. A quantum code with parameters $(\!(n,K,\delta)\!)_{2}$ is then able to encode a $K$-dimensional system into $n$ qubits, such that any errors acting on $\lfloor\tfrac{\delta-1}{2}\rfloor$ physical qubits can be corrected. One of the fundamental problems of quantum coding theory is [26, 16]:
@@
Does there exist a quantum code with given parameters $(\!(n,K,\delta)\!)_{2}$?
To date, partial answers to this questions are known in form of the quantum linear programming bounds and analytical upper bounds. Our main results are the following.
@@
Result A is shown in Section 4, Theorem 6. The proof is based on the characterization of quantum error correction through the Knill-Laflamme conditions, the quantum MacWilliams identity, and the state polynomial optimization framework to optimize over valid quantum weight enumerators. From an intermediate level of the hierarchy we recover the following for self-dual quantum codes:
@@
The Lovász number bounds the existence of self-dual quantum codes. The symmetrization over distance preserving automorphisms of a minor variation of the Lovász bound recovers the quantum Delsarte bound.
@@
Result C is shown in Section 8, Theorem 21; and Section 9, Observations 23 and 24. The proof is based on symmetrizing an intermediate level of the hierarchy over distance preserving automorphisms of the Pauli basis $\mathcal{E}_{n}$ that leave the identity invariant. The non-existence proof for the $(\!(7,1,4)\!)_{2}$ code is provided as an infeasibility certificate.
@@ 2.1. Proof sketch: SDP hierarchy
We now give a sketch of the proof of Result A. Consider a quantum channel $\mathcal{N}(\varrho)=\sum_{E\in\mathcal{N}}E\varrho E^{\dagger}$.
A quantum code requires that the Knill-Laflamme conditions [34] are met: A projector $\Pi$ corresponds to a quantum code that is able to correct errors induced by $\mathcal{N}$, if and only if
@@
In this paper we consider errors $E_{a}$ of a complete orthonormal $n$ qubit tensor-product basis ${\mathcal{E}_{n}}$, having the form $E_{a}=\sigma_{a_{1}}\otimes\dots\otimes\sigma_{a_{n}}$ where $\sigma_{0}=\mathds{1}$ and $\operatorname{tr}(\sigma_{i}^{\dagger}\sigma_{j})=\delta_{ij}$. The weight $\operatorname{wt}(E_{a})$ of an operator $E_{a}$ is the number of positions it acts non-trivially on. A code has distance $\delta$, if all errors of weight $\lfloor\tfrac{\delta-1}{2}\rfloor$ can be corrected. The size of the code space is $K=\operatorname{tr}(\Pi)$ with $\Pi$ being a projector acting on a $n$ qubit Hilbert space.
We now ask whether a code with parameters $(\!(n,K,\delta)\!)_{2}$ exists. This question can formulated as a feasibility problem:
find $\Pi$
subject to $\Pi^{2}=\Pi\,,$
$\operatorname{tr}(\Pi)=K\,,$
(2) $\Pi E_{a}^{\dagger}E_{b}\Pi=c_{ab}\Pi\quad\quad\text{for all}\quad E_{a},E_{b}\in\mathcal{E}_{n}\,:\,\operatorname{wt}(E_{a}^{\dagger}E_{b})<\delta\,,$
where $c_{ab}\in\mathds{C}$. A code with parameters $(\!(n,K,\delta)\!)_{2}$ exists, if and only if the program  2.1 is feasible.
In this work we show how problem (2.1) can be approximated by a semidefinite programming (SDP) hierarchy. The hierarchy is complete, in the sense that a set of parameters for which no code exists will be detected at some level of the hierarchy. Our proof is based on the following observations:
@@
    The Knill-Laflamme conditions for a projector $\Pi$ or rank $K$ to correspond to a quantum code are nonlinear in $\Pi$. That is, they can be formulated as the requirement that $KB_{j}(\Pi)=A_{j}(\Pi)$ for all $j=0,\dots,\delta-1$ [55], where
    (3) $A_{j}(\Pi)=\sum_{\begin{subarray}{c}E\in\mathcal{E}_{n}\\ \operatorname{wt}(E)=j\end{subarray}}\operatorname{tr}(E\Pi)\operatorname{tr}(E^{\dagger}\Pi)\,,\quad B_{j}(\Pi)=\sum_{\begin{subarray}{c}E\in\mathcal{E}_{n}\\ \operatorname{wt}(E)=j\end{subarray}}\operatorname{tr}(E\Pi E^{\dagger}\Pi)\,.$
@@
    The quantum MacWilliams identity allows to formulate this condition in terms of the $A_{j}(\Pi)$ alone [55]. Define $A(x,y)=\sum_{j=0}^{n}A_{j}(\Pi)x^{n-j}y^{j}$ and likewise for $B(x,y)$. The identity states that the $B_{j}(\Pi)$ can be expressed as linear combinations of the $A_{j}(\Pi)$,
@@
    First note that the conditions for a $n$-qubit state $\varrho$ to be proportional to a projector $\Pi$ of rank $K$ are nonlinear in its expectation values. To see this, note that swap operator exchanging the two tensor factors of $(\mathds{C}^{d})^{\otimes n}\otimes(\mathds{C}^{d})^{\otimes n}$ can be written as $(1,2)=2^{-n}\sum_{E\in\mathcal{E}_{n}}E^{\dagger}\otimes E$ where the sum is over the $n$-qubit Pauli basis $\mathcal{E}_{n}$  [18]. Decomposing cycles into elementary transpositions, a generalization of the swap trick then allows to evaluate traced powers of a state [27],
@@
    The fact that $\varrho=\frac{\Pi}{K}$ if and only if $\operatorname{tr}(\varrho^{m})=\frac{1}{K^{m-1}}$ for all $m\in\mathds{N}$ recovers the conditions for $\Pi$ to be a projector.
@@
    Second, note that the set of $n$-qubit Pauli operators can be defined algebraically: The characterization of algebras $\mathscr{C}$ over a complex field $\mathds{C}$ whose elements satisfy $A_{i}^{2}=1$ and $A_{i}A_{j}=(-1)^{\delta_{ij}}A_{j}A_{i}$ with $\delta_{ij}\in\{0,1\}$ was given by Gastineau-Hills [21]: These quasi-Clifford algebras $\mathscr{C}$ are isomorphic to the direct sum of Pauli operators acting on $m$-qubit systems,
    (6) $\mathscr{C}\simeq\bigoplus^{2^{r}}_{i=1}(\mathds{C}_{2})^{\otimes m}\,,$
    where $\mathds{C}_{2}$ is the space of complex $2\times 2$ matrices and $m,r\in\mathds{N}$.
    In combination with the constraints of the hierarchy, this restricts the Gelfand-Naimark-Segal construction of the state polynomial optimization framework to states and operators acting on the Hilbert space of $n$ qubits.
@@
    Third, the framework of state polynomial optimization [33] (also known as scalar extension [53, 54]) allows us to optimize over expressions that are nonlinear in expectation values of a state, such as $A_{j}(\varrho)=\sum_{E\in\mathcal{E}_{n},\,\operatorname{wt}(E)=j}\operatorname{tr}(E\varrho)\operatorname{tr}(E^{\dagger}\varrho)$. This is done by a variant of the Navascués-Pironio-Acín hierarchy [51] for non-commutative optimization. It consists of a hierarchy of semidefinite programs whose entries converge to monomials in the expectation values of operators on a state.
This way, the state polynomial optimization hierarchy allows us to optimize over the set of possible weight enumerators arising from $n$-qubit quantum codes. This gives a complete semidefinite programming hierarchy for the existence of qubit quantum codes with parameters $(\!(n,K,\delta)\!)_{2}$. That is, a code with given parameters exist if and only if every level of the hierarchy is feasible. We remark that semidefinite programming duality allows to find infeasibility certificates to prove that no code with the given parameters exists.
2.2. Proof sketch: Lovász and quantum Delsarte bounds
We give proof sketches for Results B and  C.

### v2

@@ Abstract.
This paper provides a semidefinite programming hierarchy based on state polynomial optimization to determine the existence of quantum codes with given parameters. The hierarchy is complete, in the sense that a $(\!(n,K,\delta)\!)_{2}$ code exists if and only if every level of the hierarchy is feasible. It is not limited to stabilizer codes and thus is applicable generally. While the machinery is formally dimension-free, we restrict it to qubit codes through quasi-Clifford algebras. We derive the quantum analog of a range of classical results: first, from an intermediate level a Lovász bound for self-dual quantum codes is recovered. Second, a symmetrization of a minor variation of this Lovász bound recovers the quantum Delsarte bound. Third, a symmetry reduction using the Terwilliger algebra leads to semidefinite programming bounds of size $O(n^{4})$. With this we give an alternative proof that there is no $(\!(7,1,4)\!)_{2}$ quantum code, and show that $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$ codes do not exist.
GM and FH were supported by the Foundation for Polish Science through TEAM-NET (POIR.04.04.00-00-17C1/18-$0$). GM was also supported by NCN grant no. 2024/53/B/ST2/04103. FH’s research was funded in whole or in part by the National Science Centre, Poland 2024/54/E/ST2/00451, the Polish National Agency for Academic Exchange under the Strategic Partnership Programme grant BNI/PST/2023/1/00013/U/00001, the Agence National de Recherche grant ANR-23-CPJ1-0012-01, and the Region Nouvelle-Aquitaine grant 34982420. AN was supported by QLCI. Much of the work was written while FH and GM were affiliated with the Jagiellonian University Kraków, Poland; while FH was affiliated with the Bordeaux Computer Science Laboratory (LaBRI), University of Bordeaux, France; and while AN was affiliated with the Duke Quantum Center, Duke University, USA. For the purpose of Open Access, the author has applied a $C$-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission. We thank Christine Bachoc, Simeon Ball, Bram Bekker, Sébastien Designolle, Dion Gijswijt, Marcus Grassl, Carlos de Gois, Otfried Gühne, Kiara Hansenne, Igor Klep, Victor Magron, Moisés Bermejo Morán, Claudio Procesi, Eric Rains, René Schwonnek, Jurij Volčič, Andreas Winter, Zhen-Peng Xu, and Karol Życzkowski for fruitful discussions and helpful comments. We especially want to thank the anonymous Referee “2” and Dion Gijswijt for extensive comments, which really helped to improve this manuscript.
@@ 1. Introduction
Coding theory studies how to store or send messages in the presence of noise. Typically, a message consists of a bit string of which a noisy channel may flip some bits. A classical error-correcting code allows to encode messages in a redundant manner, so that it can be recovered after being affected by noise. In contrast, a quantum message is represented by a positive semidefinite matrix of trace one on a $K$-dimensional Hilbert space $\mathds{C}^{K}$, that is, $\varrho\in L(\mathds{C}^{K})$ with $\varrho\succeq 0$ and $\operatorname{tr}(\varrho)=1$. A noisy quantum channel can be written in a Kraus decomposition as $\mathcal{N}(\varrho)=\sum_{E\in\mathcal{N}}E\varrho E^{\dagger}$ with $\sum_{E\in\mathcal{N}}E^{\dagger}E\preceq\mathds{1}$. A quantum error-correcting code then encodes a $K$-dimensional Hilbert space to a subspace of a $n$-qubit Hilbert space $(\mathds{C}^{2})^{\otimes n}$, such that the action of a noisy quantum channel can be corrected. The Knill-Laflamme conditions provide necessary and sufficient conditions for this to be possible [36]. Formulated in terms of the projector $\Pi$ onto the code subspace they read: if $\Pi E^{\dagger}F\Pi=c_{EF}\Pi$ for all errors $E,F$ appearing in the channel $\mathcal{N}$, then there exists a recovery map $\mathcal{R}$ such that $\mathcal{R}(\mathcal{N}(\varrho))\propto\varrho$. The size $K$ of the quantum code equals the rank of $\Pi$.
@@
A fundamental problem in both classical and quantum coding theory is to determine the maximum code size $K$ for a given block-length $n$ and minimum distance $\delta$. The maximum size can be bounded with a linear program using weight enumerators [13, 71]: for a classical linear code  $C$, the weight enumerator counts the number of codewords of Hamming weight $j$, $A_{j}(C)=|\{c\in C\,:\,\operatorname{wt}(c)=j\}|$. Here $\operatorname{wt}(c)$ is the number of non-zero positions in the bit string $c$. In contrast, the quantum weight enumerators $A_{j}(\Pi)$ decompose the Hilbert-Schmidt norm $\operatorname{tr}(\Pi^{\dagger}\Pi)$ according to the weights of quantum correlations [29]: one has $A_{j}=\sum_{E\in\mathcal{E}_{n},\,\operatorname{wt}(E)=j}\operatorname{tr}(\Pi E^{\dagger})\operatorname{tr}(\Pi E)$, where $\mathcal{E}_{n}$ is the $n$-qubit Pauli basis and $\operatorname{wt}(E)$ the number of coordinates in which $E$ acts nontrivially. The dual enumerator is given by $B_{j}=\sum_{E\in\mathcal{E}_{n},\,\operatorname{wt}(E)=j}\operatorname{tr}(\Pi E^{\dagger}\Pi E)$ 111 For stabilizer and codeword-stabilized quantum codes, the weight enumerators allow for combinatorial interpretations [51, 39], but for general quantum codes no such interpretation is known.. The introduction of weight enumerators allowed for efficiently computable linear programming bounds on the parameters $n$, $K$, and $\delta$ that a code may take. In the classical case, these are strengthened by a complete hierarchy of semidefinite programs [44, 40]. However, no such hierarchy for the quantum case exists in the literature. The aim of this paper is to fill this gap and, similar to the classical case, link it to the Lovász theta number of a graph and a Delsarte bound.
@@
Our main result is a complete semidefinite programming (SDP) hierarchy for the existence of quantum codes [Result A, Section 4]: a code with given parameters $(\!(n,K,\delta)\!)_{2}$ exist if and only if every level of the hierarchy is feasible. Inspired by [47], our hierarchy is based on the state polynomial optimization framework, which provides a sequence of outer approximations to the set of non-linear quantum correlations. This allows us to approximate the quantum weight enumerators $A_{j}$ and $B_{j}$, onto which the Knill-Laflamme and other conditions can be imposed. While this machinery is formally dimension-free, we pull this result back to $n$-qubit systems through the characterization of quasi-Clifford algebras by Gastineau-Hills [21].
We then derive quantum Lovász and Delsarte bounds. Given a graph, a subset of vertices is termed independent if no two of its vertices are adjacent. The Lovász theta number $\vartheta$ then bounds the independence number $\alpha$ of a graph, that is the largest cardinality of an independent set as $\alpha(G)\leq\vartheta(G)$. The computation of $\alpha$ is NP-hard [33], but $\vartheta$ itself can be expressed as a semidefinite program [43]. This provides an efficiently computable upper bound on $\alpha$. A classical code corresponds to a subset of the graph of all binary words of length $n$ for which words of Hamming distance less than $\delta$ are adjacent. As a consequence, the Lovász theta number also provides an upper bound on the maximum code size. A suitable symmetrization of the SDP for the Lovász theta number yields the classical Delsarte bound, which encodes the condition that the primal and dual weight enumerators are non-negative. The derivation relies on the fact that, given the indexing vector of a code $\chi$, the matrix $\chi\chi^{T}$ is positive semidefinite and has zero entries when the row and column strings have a distance in $1,\dots,\delta-1$ [22].
We show that for self-dual quantum codes 222 Self-dual codes, that is, quantum codes with $K=1$, are always assumed to be pure, meaning they must satisfy $\operatorname{tr}(E\Pi)=0$ for all $0<\operatorname{wt}(E)<\delta$. Otherwise the Knill-Laflamme conditions would be satisfied trivially., quantum Lovász and Delsarte bounds arise in a similar fashion [Result B, Section 7], with the added constraint that the matrix also has zeros in rows and columns whose Pauli strings anticommute.
Finally, following the strategy of Ref. [23] we give efficiently computable semidefinite programming bounds [Result C, Section 8]: similar to the classical case, these scale as $O(n^{4})$, where $n$ is the block length, and build on the block-diagonalization of the Terwilliger algebra of the non-binary Hamming scheme. With this, we give an an alternative proof that there is no $(\!(7,1,4)\!)_{2}$ quantum code 333 The original proof in Ref. [31] does not make use of mathematical optimization., as well as prove the non-existence of quantum codes with parameters $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$ [Section 9]. This shows that the method can lead to tighter bounds than the “enhanced” quantum linear programming bounds that include the shadow inequalities by Rains [63, 62], which do not recover our results.
@@
Overall, our approach is applicable to both stabilizer and non-stabilizer codes. It is based on moment matrices and thus offers high flexibility on the constraints to be included, so that small relaxations can be constructed. Lastly, because our formalism is dimension-free, it can in principle also target scenarios where the Hilbert space is unknown.
@@ 2. Contributions
Our aim is to determine the existence of quantum codes with block-length $n$, code distance $\delta$, and the dimension of the code space $K$. A quantum code with parameters $(\!(n,K,\delta)\!)_{2}$ is then able to encode a $K$-dimensional system into $n$ qubits, such that errors acting on at most $\lfloor\tfrac{\delta-1}{2}\rfloor$ physical qubits can be corrected. A fundamental problem in quantum coding theory is [27, 16]:
@@
Does there exist a quantum code with parameters $(\!(n,K,\delta)\!)_{2}$?
Partial answers to this questions are known in form of the quantum linear programming [71, 59] and analytical upper bounds [61, 49, 52, 30, 25]. Our main results are the following:
@@
Result A is shown in Section 4, Theorem 7. The proof is based on the characterization of quantum error correction through the Knill-Laflamme conditions, the quantum MacWilliams identity, and the state polynomial optimization framework to optimize over valid quantum weight enumerators. From an intermediate level of the hierarchy we recover the following result for self-dual quantum codes:
@@
The Lovász theta number $\vartheta$ bounds the existence of self-dual quantum codes. The symmetrization over distance preserving automorphisms of a minor variation of the Lovász bound recovers the quantum Delsarte bound.
@@
Result C is shown in Section 8, Theorem 22; and Section 9, Observations 24 and 25. The proof is based on symmetrizing an intermediate level of the hierarchy over distance preserving automorphisms of the Pauli basis $\mathcal{E}_{n}$ that leave the identity invariant. The non-existence proof for the $(\!(7,1,4)\!)_{2}$ code is provided as an infeasibility certificate in Appendix B. The certificates for $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$ codes can be found online in https://github.com/ganglesmunne/SDP_bounds_for_quantum_codes [Ref. [48]].
@@ 2.1. Proof sketch: SDP hierarchy
We now give a sketch of the proof of Result A. Consider a quantum channel $\mathcal{N}(\varrho)=\sum_{E\in\mathcal{N}}E\varrho E^{\dagger}$. A quantum code requires that the Knill-Laflamme conditions [35] are met: A projector $\Pi$ corresponds to a quantum code able to correct errors induced by $\mathcal{N}$, if and only if
@@
In this paper we consider errors that are Pauli strings, that is tensor products of Pauli matrices and the identity. The set of Pauli strings $\mathcal{E}_{n}=\{I,X,Y,Z\}^{\otimes n}$ forms a complete orthonormal basis of operators acting on $(\mathds{C}^{2})^{\otimes n}$. The weight $\operatorname{wt}(E_{a})$ of an operator $E_{a}\in\mathcal{E}_{n}$ is the number of positions it acts non-trivially on. If a code has distance $\delta$, then all errors of weight $\lfloor\tfrac{\delta-1}{2}\rfloor$ can be corrected. The size of the code space is $K=\operatorname{tr}(\Pi)$ with $\Pi$ being a projector acting on a $n$ qubit Hilbert space.
By Eq. (1), an $(\!(n,K,\delta)\!)_{2}$ code exists if and only if the following program is feasible:
$\operatorname{find}\quad$ $\Pi$
subject to $\Pi^{2}=\Pi\,,\quad\operatorname{tr}(\Pi)=K\,$,
(2) $\Pi E_{a}^{\dagger}E_{b}\Pi=c_{ab}\Pi\quad\quad\text{for all}\quad E_{a},E_{b}\in\mathcal{E}_{n}\,:\,\operatorname{wt}(E_{a}^{\dagger}E_{b})<\delta\,.$
Here $c_{ab}\in\mathds{C}$ and $\Pi$ acts on $(\mathds{C}^{2})^{\otimes n}$.
We show how the feasibility of the program (2.1) can be formulated as an SDP hierarchy. The hierarchy is complete, in the sense that an $(\!(n,K,\delta)\!)_{2}$ code exist if and only if every level of the hierarchy is feasible. It rests on the following observations:
@@
    The Knill-Laflamme conditions for a projector $\Pi$ of rank $K$ to correspond to a quantum code are nonlinear in $\Pi$. That is, they can be formulated as the requirement that $KB_{j}(\Pi)=A_{j}(\Pi)$ for all $j=0,\dots,\delta-1$ [59], where
    (3) $A_{j}(\Pi)=\sum_{\begin{subarray}{c}E_{a}\in\mathcal{E}_{n}\\ \operatorname{wt}(E_{a})=j\end{subarray}}\operatorname{tr}(E_{a}\Pi)\operatorname{tr}(E_{a}^{\dagger}\Pi)\,,\quad B_{j}(\Pi)=\sum_{\begin{subarray}{c}E_{a}\in\mathcal{E}_{n}\\ \operatorname{wt}(E_{a})=j\end{subarray}}\operatorname{tr}(E_{a}\Pi E_{a}^{\dagger}\Pi)\,.$
@@
    The quantum MacWilliams identity allows one to formulate this condition in terms of the $A_{j}(\Pi)$ alone [59]. Define $A(x,y)=\sum_{j=0}^{n}A_{j}(\Pi)x^{n-j}y^{j}$ and likewise for $B(x,y)$. The identity states that the $B_{j}(\Pi)$ can be expressed as linear combinations of the $A_{j}(\Pi)$,
@@
    First note that the conditions for a $n$-qubit state $\varrho$ to be proportional to a projector $\Pi$ of rank $K$ are nonlinear in its expectation values. To see this, note that swap operator exchanging the two tensor factors of $(\mathds{C}^{d})^{\otimes n}\otimes(\mathds{C}^{d})^{\otimes n}$ can be written as $(1,2)=2^{-n}\sum_{E_{a}\in\mathcal{E}_{n}}E_{a}^{\dagger}\otimes E_{a}$ where the sum is over the $n$-qubit Pauli basis $\mathcal{E}_{n}$  [18]. Decomposing cycles into elementary transpositions, a generalization of the swap trick then allows one to evaluate traced powers of a state [28],
@@
    The fact that $\varrho=\Pi/K$ if and only if $\operatorname{tr}(\varrho^{m})=1/K^{m-1}$ for all $m\in\mathds{N}$ recovers the conditions for $\Pi$ to be a projector.
@@
    Second, note that the set of $n$-qubit Pauli operators can be defined algebraically: The characterization of algebras $\mathscr{C}$ over a complex field $\mathds{C}$ whose elements satisfy $\alpha_{i}^{2}=1$ and $\alpha_{j}\alpha_{i}=(-1)^{\chi_{ij}}\alpha_{i}\alpha_{j}$ with $\chi_{ij}\in\{0,1\}$, $1\leq i<j\leq m$, was given by Gastineau-Hills [21]. These quasi-Clifford algebras $\mathscr{C}$ are isomorphic to a direct sum of Pauli operators acting on $s$-qubit systems,
    (6) $\mathscr{C}\simeq\bigoplus^{2^{r}}_{i=1}(\mathds{C}_{2})^{\otimes s}\,$,
    where $\mathds{C}_{2}$ is the space of complex $2\times 2$ matrices and $r+2s=m$. In combination with the constraints of the hierarchy, this restricts the Gelfand-Naimark-Segal construction of the state polynomial optimization framework to states and operators acting on the Hilbert space of $n$ qubits.
@@
    Third, the framework of state polynomial optimization [34] (also known as scalar extension [57, 58]) allows us to optimize over expressions that are nonlinear in expectation values of a state, such as $A_{j}(\varrho)=\sum_{E_{a}\in\mathcal{E}_{n},\,\operatorname{wt}(E_{a})=j}\operatorname{tr}(E_{a}\varrho)\operatorname{tr}(E_{a}^{\dagger}\varrho)$. This is done by a variant of the Navascués-Pironio-Acín hierarchy [55] for non-commutative optimization. It consists of a hierarchy of semidefinite programs whose objective value converges to that of a non-commutative state polynomial optimization problem.
This way, the state polynomial optimization hierarchy allows us to optimize over the set of possible weight enumerators arising from $n$-qubit quantum codes. This gives a complete semidefinite programming hierarchy for the existence of qubit quantum codes with parameters $(\!(n,K,\delta)\!)_{2}$. That is, a code with given parameters exist if and only if every level of the hierarchy is feasible. Semidefinite programming duality can then be used to find infeasibility certificates to prove that no code with the given parameters exists.
2.2. Proof sketch: Lovász, Delsarte, and symmetry-reduced SDP bounds
We sketch the proofs of Results B and  C.